OpenStudy (anonymous):

what is the sum of 12 and a number is 4 times the number as written as an equation?

OpenStudy (anonymous):

sum means add, so if the number was \(n\), how would you write the sum of n and 12?

OpenStudy (anonymous):

12n?

OpenStudy (anonymous):

12n is how you would multiply. The sum of 12 and n means "12 plus n"

OpenStudy (anonymous):

My options are 12+4n = n 12+n=4n 12n=4+n 12(n+4)=n

OpenStudy (anonymous):

So the answer would be C?

OpenStudy (anonymous):

no. look at what I just typed above about multiplication/addition

OpenStudy (anonymous):

oooh i see

OpenStudy (anonymous):

Sorry I didn't see your response

OpenStudy (anonymous):

np.

OpenStudy (anonymous):

it would be B right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

you're welcome